### v2

# LeTO: Learning Constrained Visuomotor Policy with Differentiable Trajectory Optimization

Abstract
This paper introduces LeTO, a method for learning constrained visuomotor policy via differentiable trajectory optimization. Our approach integrates a differentiable optimization layer into the neural network. By formulating the optimization layer as a trajectory optimization problem, we enable the model to end-to-end generate actions in a safe and constraint-controlled fashion without extra modules. Our method allows for the introduction of constraints information during the training process, thereby balancing the training objectives of satisfying constraints, smoothing the trajectories, and minimizing errors with demonstrations. This “gray box” method marries the optimization-based safety and interpretability with the powerful representational abilities of neural networks. We quantitatively evaluate LeTO in simulation and on the real robot. In simulation, LeTO achieves a success rate comparable to state-of-the-art imitation learning methods, but the generated trajectories are of less uncertainty, higher quality, and smoother. In real-world experiments, we deployed LeTO to handle constraints-critical tasks. The results show the effectiveness of LeTO comparing with state-of-the-art imitation learning approaches. We release our code at https://github.com/ZhengtongXu/LeTO.
Index Terms:
Robotic manipulation, imitation learning, differentiable optimization.
I Introduction
Imitation learning [1] focuses on deriving robot policies from demonstrations. This process can be formulated as a supervised learning task, aiming to learn the mapping between observations and robot actions. Through imitation learning, robots can perform highly complex and diverse tasks [2, 3, 4, 5, 6, 7, 8].
However, ensuring the safety of robots is very important due to the interaction between robots and the real world. In contrast to optimization-based trajectory generation methods, imitation learning often exhibits greater uncertainty, which can lead to system instability, reduced robustness, and safety concerns. Hence, in comparison to other supervised learning problems, this pose novel and exceptional challenges in safety for imitation learning.
Existing works focus on different aspects and challenges of imitation learning, such as improving action accuracy by addressing compounding errors [4] and represent multi-modal distributions by implicit policy [9, 10, 2]. However, these methods overlook the critical aspects of safety and smoothness of generated trajectories. For robotic systems, their behaviors need to meet specific constraints to ensure the safety of the system. Moreover, for certain tasks, the robot actions also need to meet specific constraints to ensure their successful completion. In addition, adding explainable constraints for neural network-based algorithms, often called “black boxes”, is much harder than for traditional model-based methods.
[…]
Figure 1: In LeTO, we enable the model to end-to-end generate actions in a safe and constraint-controlled fashion without extra modules. To the best of our knowledge, LeTO is the first visuomotor imitation learning framework that not only utilizes differentiable optimization but also demonstrates its efficacy in real-world robotic manipulation tasks.
In this paper, we propose LeTO, learning constrained visuomotor policy with differentiable trajectory optimization. The strengths and novelties of our method are as follows.
1) Differentiable optimization layer: We integrate a differentiable optimization layer to the neural network and formulate it as a trajectory optimization problem. We demonstrate that the differentiable optimization layer we design has optimization feasibility during training and model representation capability. To the best of our knowledge, our work is the first visuomotor imitation learning framework that not only utilizes differentiable optimization but also demonstrates its efficacy in both simulation and real-world robotic manipulation tasks.
2) Safe and constrained action generation: In our approach, we formulate the optimization layer as a trajectory optimization problem. Through this approach, the policy generate constrained trajectories, enhancing the overall safety and smoothness of robot actions. Since the model incorporates position, velocity, and acceleration constraints during end-to-end training, it ensures robot safety without compromising performance when deploying the model as a real-time policy. Compared to the “black box” characteristics of neural networks, our approach can be described as a “gray box” that combines the safety and interpretability of optimization-based trajectory generation with the powerful representational capabilities of neural networks.
We evaluate LeTO both in simulations (Section IV) and on real robot tasks (Section V) quantitatively, and benchmark with diffusion policy [11], LSTM-GMM [12], and IBC [10]. In the simulated tests, LeTO achieves a success rate comparable to diffusion policy and better than LSTM-GMM and IBC, but the generated trajectories are of less uncertainty, higher quality, and smoother comparing with these methods. On the real robot, we implement constraints-critical manipulation tasks using LeTO, which highlights LeTO’s strengths.
II Related work
II-A Trajectory Optimization
Trajectory optimization is pivotal in generating safe and smooth trajectories for robots and remains an active area of research. CHOMP [13] approaches trajectory optimization by gradient techniques, focusing on optimizing the trade-off between safety and smoothness. TrajOpt [14] uses a sequential convex optimization to generate smooth and collision-free trajecotries and can use naive straight-line initializations that might be in collision. The minimum-snap trajectory optimization [15] uses piecewise polynomials to represent the trajectory and is optimized by quadratic programs. The work in [16] proposes a method for reformulating nondifferentiable collision avoidance constraints into smooth, differentiable constraints and enables real-time optimization-based trajectory planning.
In LeTO, we innovatively combine trajectory optimization with imitation learning by differentiable optimization, to generate smooth and constraint-compliant trajectories during policy rollout. Our model is trained end-to-end. During the training process, LeTO not only learns the policy from human demonstrations but also learns the appropriate trajectory optimization parameters suitable for the policy.
II-B Imitation Learning
The basic way of imitation learning explicitly maps observations to actions [1, 17, 18, 19, 20, 21]. It can be trained using a regression loss. However, these policies are not ideal for capturing multi-modal distributions [10].
Prior works have aimed to represent multi-modal distributions by converting the regression into classification [22, 23, 24], using action discretization coupled with a multi-task action correction [25], using MDNs [12], and using implicit modeling that including energy-based model [10, 9] and diffusion model [2]. However, these methods overlook the critical aspects of safety for robot systems. For robotic systems, their behaviors need to meet specific constraints to ensure the safety of the system and the successful completion of tasks.
LeTO focuses on combining explainable and model-based trajectory optimization with imitation learning. In this way, LeTO can be end-to-end trained and generate action in a safe and constraint-controlled fashion without extra modules.
II-C Differentiable Optimization in Robot Learning
Previous works utilize differentiable optimization for modeling discontinuous functions and dynamics [26, 27]. Additionally, some works focus on combining model-based methods with neural networks through differentiable optimization, such as using differentiable MPC [28] and koopman operator [29]. However, these methods are not suitable for tasks with high-dimensional observations, such as using camera images from multiple viewpoints.
The work in [30] proposes a tactile-reactive grasping controller that combines image encoder and differentiable MPC. However, it can not be used for more general policy learning.
For robot navigation and obstacle avoidance, various end-to-end learning frameworks embed with differentiable optimization are proposed, such as using control barrier function [31], gradient-based correction [32], and stacking prediction, planning, and control in a differentiable way [33]. However, all these methods are not suited for performing manipulation tasks.
DiffTOP [34] is a method that integrate differentiable trajectory optimization as the policy representation to generate actions. In integrating imitation learning, DiffTOP focuses on capturing multi-modal distributions. However, it can not generating safe and constrained trajectories. In contrast, LeTO focuses on how the integration of differentiable trajectory optimization into the imitation learning framework enables the generation of trajectories that not only perform tasks effectively but also satisfy constraints.
[…]
Figure 2: Overview of LeTO. We enable the model to end-to-end generate actions in a safe and constraint-controlled fashion by integrating a differentiable trajectory optimization layer.
Another category of methods closely related to LeTO is riemannian motion policy (RMP) [35] and its extensions [36, 37]. Specifically, the works in [36, 37] enhance RMP’s integration into end-to-end robot learning by incorporating automatic differentiation. This enables end-to-end training and inference for policies with RMP structures. Methods based on RMP are particularly focused on generating motion in spaces with high degrees of freedom and nonlinear dynamics. In contrast, LeTO offers a capable form of generating constrained motion at the task space trajectory level with differentiable optimization. This advantage makes it better suited for adaptation to visuomotor policies and demonstrates superior performance in real tasks. To the best of our knowledge, our work is the first visuomotor imitation learning framework that not only utilizes differentiable optimization but also demonstrates its efficacy in both simulation and real-world robotic manipulation tasks.
[…]
Figure 3: Illustration of training data sampling.
III Approach
In this paper, we assume access to an offline trajectories dataset of the task we want to do. The goal is to learn a policy from this dataset offline and can successfully do the task by running the policy online. The overview of LeTO is shown in Fig. 2. In this section, we will detail the design of the model.
III-A Training Data Sampling
In this paper, we categorize actions into two types: one is discrete actions, such as grasping and releasing; the other is continuous actions, such as the motion of the robot. For continuous actions, we consider the velocity/increment of the end-effector as the action [12, 38, 39, 25, 17, 18]. Denote the dataset as
$\mathcal{D}=\left\{\left(o_{0}^{i},y_{0}^{i},o_{1}^{i},y_{1}^{i},\ldots,o_{T^{i}_{d}}^{i},y_{T^{i}_{d}}^{i}\right)\right\}_{i=1}^{N_{d}},$
where $o$ represents the observations of the robot, such as images of the cameras with different views and robot states and $y\in\mathbb{R}^{D_{y}}$ represents the demonstrated action aligned with the observation $o$. $N_{d}$ is the total number of the demonstrated trajectories, $T_{d}$ is the total length of a trajectory, and $D_{y}$ is the dimension of action $y$.
Inspired by model predictive control and diffusion policy [2], in LeTO, we predict a action sequence at every time step. Moreover, we care about how to predict trajectories that satisfy constraints rather than action sequences simply fitting the human demonstration. Therefore, before training, we first sample many fixed-length sequences in following format from the dataset
$\left(o_{t_{0}},\mathbf{y}_{t_{0}},o_{{t_{0}}+1},\mathbf{y}_{{t_{0}}+1},\ldots,o_{t_{0}+T_{s}-1},\mathbf{y}_{t_{0}+T_{s}-1}\right),$
where $\mathbf{y}_{t}=\left[{y}_{t}^{\mathrm{T}},{y}_{t+1}^{\mathrm{T}},\ldots,{y}_{t+T_{p}-1}^{\mathrm{T}}\right]^{\mathrm{T}}\in\mathbb{R}^{T_{p}D_{y}}$. $T_{s}$ is the length of each data sampling and $T_{p}$ is the length of the predicted action sequence, as shown in Fig. 3. The learning objective is minimizing the errors between the predicted $\hat{\mathbf{y}}^{*}_{t}$ and $\mathbf{y}_{t}$ (e.g. using MSE loss to assess the distance between them), as shown in Fig. 2. During policy rollout, the robot will execute the first $T_{a}$-step actions of $\hat{\mathbf{y}}^{*}_{t}$, and then repeat the same process based on new observations. This process is typically called receding horizon planning/control.
Continuous actions need to be constrained to ensure actions safety and smoothness while discrete actions do not. Therefore, we define a selection matrix $S$ to pick the continuous action vector $v$ that require to be constrained.
$v=Sy\in\mathbb{R}^{D_{v}}.$ (1)
As mentioned earlier, we focus on a velocity-based action space, so here, $v$ refers to the velocity at which the robot is being controlled and $D_{v}$ is the dimension of $v$.
III-B Observation Encoder
The observation encoder maps the observations, like raw RGB images and robot states, into a latent embedding $e$ and is trained end-to-end with LeTO. To capture the sequential correlation of the task, we use a recurrent neural network (RNN) architecture, $E_{\theta}$, where the right subscript $\theta$ represent the parameters of the network. We choose long short-term memory (LSTM) as our RNN model. At each time step, $e_{t},h_{t+1}=E_{\theta}\left(o_{t},h_{t}\right),$ where $h$ is the hidden state. We define $e_{t}\in\mathbb{R}^{T_{p}D_{y}}$. More detailed reasons of setting the dimension as $T_{p}D_{y}$ can be seen at Section III-C.
III-C Differentiable Trajectory Optimization Layer
The differentiable trajectory optimization (DTO) layer maps the embedding $e_{t}$ to a predicted action sequence $\hat{\mathbf{y}}_{t}=\left[\hat{y}_{t}^{\mathrm{T}},\hat{y}_{t+1}^{\mathrm{T}},\ldots,\hat{y}_{t+T_{p}-1}^{\mathrm{T}}\right]^{\mathrm{T}}\in\mathbb{R}^{T_{p}D_{y}}$. In this section, we will show how to design this layer based on trajectory optimization, learn this layer during training, and integrate it with the whole neural network.
By equation (1), the predicted velocity command sequence
$\hat{\mathbf{v}}_{t}=\mathbf{S}\hat{\mathbf{y}}_{t}=\left[\hat{v}_{t}^{\mathrm{T}},\hat{v}_{t+1}^{\mathrm{T}},\ldots,\hat{v}_{t+T_{p}-1}^{\mathrm{T}}\right]^{\mathrm{T}}\in\mathbb{R}^{T_{p}D_{v}},$ (2)
where $\mathbf{S}=\text{blkdiag}\left(S,\dots,S\right)\in\mathbb{R}^{T_{p}D_{v}\times T_{p}D_{y}}$. Then the predicted acceleration sequence and position sequence are
$\hat{\mathbf{a}}_{t}$ $=\left[\hat{a}_{t}^{\mathrm{T}},\hat{a}_{t+1}^{\mathrm{T}},\ldots,\hat{a}_{t+T_{p}-2}^{\mathrm{T}}\right]^{\mathrm{T}}$
$=\frac{1}{\Delta t}\left[\hat{v}_{t+1}^{\mathrm{T}}-\hat{v}_{t}^{\mathrm{T}},\ldots,\hat{v}_{t+T_{p}-1}^{\mathrm{T}}-\hat{v}_{t+T_{p}-2}^{\mathrm{T}}\right]^{\mathrm{T}}$
$=\frac{1}{\Delta t}\left[\begin{matrix}-1&1&&&\\ &-1&1&&\\ &&\ddots&\ddots&\\ &&&-1&1\end{matrix}\right]\hat{\mathbf{v}}_{t}$ (7)
$=\mathbf{A}_{\text{diff}}\hat{\mathbf{v}}_{t}\in\mathbb{R}^{T_{p}D_{v}-D_{v}},$ (8)
$\hat{\mathbf{p}}_{t}$ $=\left[\hat{p}_{t+1},\ldots,\hat{p}_{t+T_{p}-1}\right]^{\mathrm{T}}$
$=\left[\begin{matrix}p_{t}\\ p_{t}\\ \vdots\\ p_{t}\end{matrix}\right]+\left[\begin{matrix}\Delta t&&&\\ \Delta t&\Delta t&\\ \vdots&\vdots&\ddots&\\ \Delta t&\Delta t&\ldots&\Delta t\end{matrix}\right]\hat{\mathbf{v}}_{t}$ (17)
$=\mathbf{p}_{t,0}+\mathbf{A}_{\text{inte}}\hat{\mathbf{v}}_{t}\in\mathbb{R}^{T_{p}D_{v}-D_{v}},$ (18)
where $\hat{p}_{t+i}={p}_{t}+\hat{v}_{t}\Delta t+\dots+\hat{v}_{t+i-1}\Delta t,i=1,\ldots,T_{p}-1$, and $p_{t}$ is the robot end-effector position at time step $t$.
The sequence $\hat{\mathbf{p}}_{t},\hat{\mathbf{v}}_{t},\text{and}~{}\hat{\mathbf{a}}_{t}$ together constitute the trajectory generated by the policy. To optimize the trajectory, we must consider the relationship between $\hat{\mathbf{p}}_{t},\hat{\mathbf{v}}_{t},\text{and}~{}\hat{\mathbf{a}}_{t}$ and human demonstration, as well as the constraints of $\hat{\mathbf{p}}_{t},\hat{\mathbf{v}}_{t},\text{and}~{}\hat{\mathbf{a}}_{t}$. Based on this, the forward pass of the DTO layer is the following optimization problem:
$\hat{\mathbf{y}}^{\star}_{t}$ $=\underset{\hat{\mathbf{y}}_{t}}{\operatorname{argmin}}\frac{1}{2}\hat{\mathbf{y}}_{t}^{\mathrm{T}}\mathbf{Q}\hat{\mathbf{y}}_{t}+e_{t}^{\mathrm{T}}\hat{\mathbf{y}}_{t}+\frac{\alpha}{2}\hat{\mathbf{a}}_{t}^{\mathrm{T}}\hat{\mathbf{a}}_{t},$ (19)
subject to $b_{\text{min}}\leq A_{\text{pos}}\hat{p}_{t+i}\leq b_{\text{max}},i=1,\ldots,T_{p}-1,$
$v_{\text{min}}\leq\hat{v}_{t+j}\leq{v}_{\text{max}},j=0,\ldots,T_{p}-1,$
$a_{\text{min}}\leq\hat{a}_{t+k}\leq a_{\text{max}},k=0,\ldots T_{p}-2.$
Rewrite the optimization problem (19) by equations (2), (8), and (18), then we have
$\hat{\mathbf{y}}^{\star}_{t}$ $=\underset{\hat{\mathbf{y}}_{t}}{\operatorname{argmin}}\frac{1}{2}\hat{\mathbf{y}}_{t}^{\mathrm{T}}\bar{\mathbf{Q}}\hat{\mathbf{y}}_{t}+e_{t}^{\mathrm{T}}\hat{\mathbf{y}}_{t},$ (20)
subject to
$\mathbf{b}_{\text{min}}\leq\mathbf{A}_{\text{pos}}(\mathbf{p}_{t,0}+\mathbf{A}_{\text{inte}}\mathbf{S}\hat{\mathbf{y}}_{t})\leq\mathbf{b}_{\text{max}},$ (21)
$\mathbf{v}_{\text{min}}\leq\mathbf{S}\hat{\mathbf{y}}_{t}\leq\mathbf{v}_{\text{max}},$ (22)
$\mathbf{a}_{\text{min}}\leq\mathbf{A}_{\text{diff}}\mathbf{S}\hat{\mathbf{y}}_{t}\leq\mathbf{a}_{\text{max}},$ (23)
where $~{}\bar{\mathbf{Q}}=(\mathbf{Q}+\alpha\mathbf{S}^{\mathrm{T}}\mathbf{A}_{\text{diff}}^{\mathrm{T}}\mathbf{A}_{\text{diff}}\mathbf{S}),$
$~{}\mathbf{A}_{\text{pos}}=\text{blkdiag}\left(A_{\text{pos}},\dots,A_{\text{pos}}\right)$
$~{}~{}~{}~{}~{}~{}~{}~{}\in\mathbb{R}^{(T_{p}D_{v}-D_{v})\times(T_{p}D_{v}-D_{v})}.$
The coefficient $\alpha$ controls the smoothness of the trajectory. It can be observed in (19) that the larger the value of $\alpha$, the greater the cost associated with the sum of accelerations, leading the optimization problem to generate trajectories with a smaller sum of accelerations. $A_{\text{pos}}$ and $b_{\text{max,min}}$ define the convex constraints that the position must satisfy. $v_{\text{max,min}}$ and $a_{\text{max,min}}$ define the constraints for the velocity and acceleration of the trajectory. $\alpha,A_{\text{pos}},b_{\text{max,min}},v_{\text{max,min}}$ and $a_{\text{max,min}}$ are interpretable and model-based, so we do not need to learn these parameters during training. Instead, it is advisable to specify them according to the requirements of the task. On the contrary, $\mathbf{Q}\in\mathbb{R}^{T_{p}D_{y}}$ is a parameters that we need to learn during training. Next, we will demonstrate how to learn $\mathbf{Q}$ and its specific meaning in our trajectory optimization problem.
Optimization problem (20) is a quadratic program (QP). Therefore, the layer (20) can perform forward pass and backpropagation in batch form as long as it is always feasible in the training process [40]. The feasibility of (20) can be guaranteed by ensuring $\mathbf{Q}$ symmetric positive definite [41]. To achieve that, we utilize a Cholesky factorization
$\mathbf{Q}=\mathbf{L}\mathbf{L}^{\mathrm{T}}+\epsilon\mathbf{I},$
and directly learn $\mathbf{L}$ , where $\mathbf{L}$ is a lower triangular matrix and $\epsilon$ is a very small scalar (e.g. $1\times 10^{-4}$). By observing optimization problem (20), Remark 1 can be derived.
Remark 1 (feasibility of the differentiable optimization layer).
If the constraint $\mathbf{b}_{\rm{min}}\leq\mathbf{A}_{\rm{pos}}\mathbf{p}_{t,0}\leq\mathbf{b}_{\rm{max}}$ is satisfied at each time step $t$, the optimization problem is always feasible, no matter how $\mathbf{L}$ and $e_{t}$ change. This means the optimization problem remains consistently feasible, which ensuring stable training process. $\mathbf{b}_{\rm{min}}\leq\mathbf{A}_{\rm{pos}}\mathbf{p}_{t,0}\leq\mathbf{b}_{\rm{max}}$ represents that the human demonstrations should satisfy the position constraints. Additionally, the human demonstrations do not need to satisfy the velocity and acceleration constraints strictly, because they have nothing to do with the feasibility of the optimization problem. Further more, for human demonstration, satisfying position constraints is much easier than satisfying the velocity and acceleration constraints.
Based on that $\mathbf{Q}$ is symmetric positive definite, the optimization problem (20) can be converted as the following least squares problem:
$\hat{\mathbf{y}}^{\star}_{t}=\underset{\hat{\mathbf{v}}_{t}}{\operatorname{argmin}}\frac{1}{2}\|\bar{\mathbf{L}}^{\mathrm{T}}\hat{\mathbf{y}}_{t}-{\bar{e}_{t}}\|^{2},$ (24)
[…]
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}{e_{t}}=-\bar{\mathbf{L}}{\bar{e}_{t}}.$
Based on Cholesky factorization, $\bar{\mathbf{L}}$ is a lower triangular matrix with real and positive diagonal entries. By observing optimization problem (24), Remark 2 can be derived.
Remark 2 (representational power).
DTO layer actually optimize a linear transformation under motion constraints. Since $\bar{\mathbf{L}}$ is a lower triangular matrix and it is fully ranked, the linear transformation here is
$\hat{\mathbf{y}}^{\star}_{t}=-({\bar{\mathbf{L}}^{\mathrm{T}}})^{-1}\bar{\mathbf{L}}^{-1}e_{t}.$
Since the observation encoder is just a normal LSTM architecture, it is already has the power of representing the policy. Therefore, the whole model including the DTO layer has the power of fitting the human demonstration.
III-D Policy Deployment
When the policy is deployed on the robot, the RNN is unrolled one-step at a time, $e_{t},h_{t+1}=E_{\theta}\left(o_{t},h_{t}\right)$. The hidden state $h$ is refreshed every $T_{s}$ steps. In addition, the training is on discrete trajectory clips while the robot trajectory is continuous in reality. Therefore, the forward pass of the DTO layer need to add a new constraint during test time to fully constrain the motion generation.
$\hat{\mathbf{y}}^{\star}_{t}=\underset{\hat{\mathbf{y}}_{t}}{\operatorname{argmin}}$ $\frac{1}{2}\hat{\mathbf{y}}_{t}^{\mathrm{T}}\bar{Q}\hat{\mathbf{y}}_{t}+e_{t}^{\mathrm{T}}\hat{\mathbf{y}}_{t},$
[…]
$a_{\text{min}}\Delta t$ $\leq\hat{v}_{t}-\hat{v}_{t-1}\leq a_{\text{max}}\Delta t$
where $\hat{v}_{t-1}$ represents the last executed action for the robot at time step $t$ (executed at time step $t-1$). Adding this new constraint is transferring the model trained on a discrete set of trajectories into continuous online trajectory optimization.
TABLE I: Summary of success rates. We use the imitation learning benchmark, RoboMimic (Visual Policy) [12]. We present the average of the maximum success rate (%) across 3 training seeds and 25 different initial environmental conditions (totaling 75 conditions). Results for IBC were sourced from [2] to enable a comparison with conventional methods. The black bold font indicates the highest success rate among all tasks. Notably, for the square task, LeTO achieves a comparable success rate to diffusion policy (bold in green) and surpasses diffusion policy with constraints clipping.
Can Square
$\mathrm{ph}$ $\mathrm{mh}$ $\mathrm{ph}$ $\mathrm{mh}$
IBC [10] 8 0 3 0
LSTM-GMM [12] $94.6\pm 2.3$ $94.6\pm 2.3$ $78.7\pm 6.1$ $77.3\pm 8.3$
LSTM-GMM clipping $90.7\pm 4.6$ $89.3\pm 2.3$ $73.3\pm 8.3$ $70.7\pm 2.3$
DiffusionPolicy [2] $\textbf{100.0}\pm\textbf{0.0}$ $\textbf{100.0}\pm\textbf{0.0}$ $\textbf{98.7}\pm\textbf{2.3}$ $\textbf{90.7}\pm\textbf{4.6}$
DiffusionPolicy clipping $98.7\pm 2.3$ $93.3\pm 4.6$ $92.0\pm 4.0$ $82.7\pm 2.3$
LeTO $\textbf{100.0}\pm\textbf{ 0.0}$ $\textbf{100.0}\pm\textbf{0.0}$ $\textbf{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@rgb@stroke{0}{1}{0}\pgfsys@color@rgb@fill{0}{1}{0}94.7}}{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@rgb@stroke{0}{1}{0}\pgfsys@color@rgb@fill{0}{1}{0}\pm}\textbf{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@rgb@stroke{0}{1}{0}\pgfsys@color@rgb@fill{0}{1}{0}4.6}}$ $\textbf{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@rgb@stroke{0}{1}{0}\pgfsys@color@rgb@fill{0}{1}{0}88.0}}{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@rgb@stroke{0}{1}{0}\pgfsys@color@rgb@fill{0}{1}{0}\pm}\textbf{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@rgb@stroke{0}{1}{0}\pgfsys@color@rgb@fill{0}{1}{0}4.0}}$
[…]
Figure 4: Simulation benchmarks: pick and place the can (can task) and grasping and assembling the square nut (square task).
IV Simulation Evaluation
We employed two simulation tasks used in [12] for our simulation evaluation: pick and place the can (can task) and grasping and assembling the square nut (square task), as shown in Fig. 4. The results demonstrate that our approach not only achieves success rates comparable to diffusion policy [2], but also significantly surpasses IBC [10] and LSTM+GMM [12]. Furthermore, owing to the presence of constraints and differentiable trajectory optimization, our method exhibits superior performance in terms of generated trajectory quality compared to other methods, which is very important for robot systems.
For training parameters of diffusion policy and LSTM-GMM, we use the default configurations provided in their released code. We also do the same for our realworld evaluation. We have released our code, training data, and checkpoints for reproducing our results.
The can and square tasks are from RoboMimic benchmark [12]. Apart from the most basic lifting task, we opted not to include the tool hanging and transport tasks. As highlighted in [40], solving optimization problems precisely, as we are doing here, exhibits cubic complexity concerning the number of variables and/or constraints. Consequently, our method’s training speed for tasks like tool hanging and transport, which involve large datasets or high dimension outputs, is sluggish, and under limited GPU computational resources, obtaining results within a short time is unfeasible. Nevertheless, as we have already demonstrated the capability of our approach to achieve high successful rates and generate high quality trajectories through the outcomes of the two tasks mentioned above, we decided against further training on the tool hanging and transport tasks.
We train policies on 3 seeds (42, 142, 242), and test in 25 different initial environments for each seeds (totaling 75 conditions). For each task, we train policies on two different types of datasets: Proficient-Human (ph), and Multi-Human (mh) datasets [12]. We report the success rate (see Table I) of all these experiments. We also report metrics for evaluating the quality of the generated trajectories (see Tables II, III, IV, and V).
TABLE II: Trajectory metrics ($10^{-2}$) of the can-ph task. The results correspond to the results in Table I.
avg-mean-lin avg-max-lin avg-std-lin avg-mean-rot avg-max-rot avg-std-rot num-of-opt
DiffusionPolicy [2] $8.23\pm 0.70$ $69.78\pm 2.67$ $8.39\pm 0.58$ $\textbf{1.03}\pm\textbf{ 0.12}$ $12.73\pm 0.78$ $\textbf{1.29}\pm\textbf{0.12}$ 2
LSTM-GMM [12] $9.78\pm 3.54$ $77.49\pm 15.16$ $10.61\pm 3.48$ $\textbf{1.04}\pm\textbf{0.27}$ $24.58\pm 9.73$ $2.71\pm 0.26$ 1
DiffusionPolicy clipping $7.73\pm 0.10$ $\textbf{10.00}\pm\textbf{0.00}$ $\textbf{3.21}\pm\textbf{ 0.06}$ $3.61\pm 0.47$ $\textbf{9.94}\pm\textbf{0.07}$ $2.45\pm 0.08$ 3
LSTM-GMM clipping $5.61\pm 1.03$ $\textbf{10.00}\pm\textbf{0.00}$ $3.68\pm 0.10$ $1.16\pm 0.13$ $\textbf{9.59}\pm\textbf{0.14}$ $\textbf{1.37}\pm\textbf{0.05}$ 3
LeTO $\textbf{4.36}\pm\textbf{0.10}$ $\textbf{10.00}\pm\textbf{0.00}$ $\textbf{3.47}\pm\textbf{0.03}$ $1.50\pm 0.21$ $\textbf{9.78}\pm\textbf{0.37}$ $1.73\pm 0.30$ 4
TABLE III: Trajectory metrics ($10^{-2}$) of the can-mh task. The results correspond to the results in Table I.
avg-mean-lin avg-max-lin avg-std-lin avg-mean-rot avg-max-rot avg-std-rot num-of-opt
DiffusionPolicy [2] $5.97\pm 0.93$ $64.85\pm 9.04$ $7.32\pm 0.85$ $\textbf{0.60}\pm\textbf{0.06}$ $10.95\pm 1.16$ $\textbf{0.96}\pm\textbf{0.09}$ 2
LSTM-GMM [12] $6.02\pm 0.78$ $72.16\pm 1.94$ $7.41\pm 0.36$ $0.75\pm 0.08$ $21.48\pm 9.64$ $1.57\pm 0.50$ 0
DiffusionPolicy clipping $7.35\pm 0.14$ $\textbf{10.00}\pm\textbf{0.00}$ $3.30\pm 0.03$ $2.74\pm 0.19$ $9.80\pm 0.07$ $2.26\pm 0.12$ 1
LSTM-GMM clipping $4.74\pm 0.46$ $\textbf{10.00}\pm\textbf{0.00}$ $3.45\pm 0.10$ $0.79\pm 0.10$ $9.16\pm 0.87$ $1.19\pm 0.43$ 1
LeTO $\textbf{2.66}\pm\textbf{0.68}$ $\textbf{10.00}\pm\textbf{0.00}$ $\textbf{2.60}\pm\textbf{0.31}$ ${0.74}\pm{0.22}$ $\textbf{8.28}\pm\textbf{1.04}$ $\textbf{0.98}\pm\textbf{0.24}$ 5
TABLE IV: Trajectory metrics ($10^{-2}$) of the square-ph task. The results correspond to the results in Table I.
avg-mean-lin avg-max-lin avg-std-lin avg-mean-rot avg-max-rot avg-std-rot num-of-opt
DiffusionPolicy [2] $8.17\pm 0.78$ $70.66\pm 4.49$ $8.65\pm 0.40$ $\textbf{0.95}\pm\textbf{0.11}$ $13.22\pm 0.80$ $\textbf{1.32}\pm\textbf{0.11}$ 2
LSTM-GMM [12] $8.19\pm 2.49$ $68.75\pm 6.52$ $8.76\pm 1.82$ $\textbf{0.91}\pm\textbf{0.08}$ $13.81\pm 3.51$ $\textbf{1.30}\pm\textbf{0.24}$ 2
DiffusionPolicy clipping $6.90\pm 0.75$ $\textbf{10.00}\pm\textbf{0.00}$ $\textbf{3.36}\pm\textbf{0.06}$ $2.09\pm 0.88$ $\textbf{9.09}\pm\textbf{0.82}$ $1.75\pm 0.59$ 3
LSTM-GMM clipping $\textbf{5.11}\pm\textbf{1.46}$ $\textbf{10.00}\pm\textbf{0.00}$ $\textbf{3.36}\pm\textbf{0.20}$ $\textbf{0.98}\pm\textbf{0.22}$ $\textbf{9.04}\pm\textbf{0.54}$ $\textbf{1.27}\pm\textbf{0.29}$ 6
LeTO $\textbf{5.08}\pm\textbf{0.33}$ $\textbf{10.00}\pm\textbf{0.00}$ $\textbf{3.35}\pm\textbf{0.03}$ $1.28\pm 0.17$ $\textbf{9.24}\pm\textbf{0.15}$ $\textbf{1.37}\pm\textbf{0.08}$ 5
TABLE V: Trajectory metrics ($10^{-2}$) of the square-mh task. The results correspond to the results in Table I.
avg-mean-lin avg-max-lin avg-std-lin avg-mean-rot avg-max-rot avg-std-rot num-of-opt
DiffusionPolicy [2] $5.42\pm 0.30$ $60.97\pm 4.38$ $6.61\pm 0.44$ $\textbf{0.57}\pm\textbf{0.08}$ $13.41\pm 3.45$ $1.11\pm 0.26$ 1
LSTM-GMM [12] $5.37\pm 0.66$ $54.43\pm 6.99$ $6.28\pm 1.00$ $0.64\pm 0.08$ $13.31\pm 2.66$ $1.04\pm 0.22$ 0
DiffusionPolicy clipping $7.04\pm 0.42$ $\textbf{10.00}\pm\textbf{0.00}$ $3.35\pm 0.06$ $2.41\pm 0.50$ $9.46\pm 0.34$ $2.04\pm 0.36$ 1
LSTM-GMM clipping $3.81\pm 0.44$ $\textbf{10.00}\pm\textbf{0.00}$ $3.07\pm 0.13$ $\textbf{0.63}\pm\textbf{0.04}$ $\textbf{7.80}\pm\textbf{0.23}$ $\textbf{0.85}\pm\textbf{0.04}$ 4
LeTO $\textbf{3.08}\pm\textbf{0.04}$ $\textbf{10.00}\pm\textbf{0.00}$ $\textbf{2.70}\pm\textbf{0.01}$ $0.91\pm 0.06$ $\textbf{8.18}\pm\textbf{0.07}$ $1.09\pm 0.07$ 4
The velocity action outputs in RoboMimic dataset [12] are normalized to [-1, 1]. Therefore, for our implementation of LeTO, we impose a normalized velocity constraint [-1,1], and we introduce an acceleration constraint [-0.1,0.1]. For the simulation tests, we do not apply any position constraints, as our aim is to solely compare the success rates and the smoothness of the trajectories generated by different methods.
For the assessment of trajectory quality, we calculate the maximum acceleration, average acceleration, and acceleration standard deviation for each dimension (x,y,z, roll, pitch, and yaw) and for each environment condition (25 environments for each seed). Then we calculate the average among dimensions and environments. For example
[…]
where $i$ represents the $i$-th environment and $a$ represents acceleration. “lin” here represents linear motion and “rot” represents rotational motion. The remaining matrics can be computed by analogy. Additionally, it is worth noting that the values of the acceleration here are obtained by computing the difference of the normalized velocity. It is directly proportional to the actual acceleration, and this proportionality constant is a fixed value for all experiments. We mark the values within the range of 100% to 110% of the minimum value in each metric as the optimal indicators (bolded), and denote the number of optimal indicators for each method as “num-of-opt”.
In order to investigate the impact of direct action clipping on success rate and trajectory smoothness for baselines, we implement a clipping method. It involves evaluating the difference between the generated action and the previously executed action to determine if it falls within the range of -0.1 to 0.1, which is the same constraints we implement for LeTO. If it exceeds this range, we apply clipping to bring it within this range. This approach constrains acceleration.
We observe that LeTO significantly enhances trajectory smoothness by trajectory optimization while maintaining a consistently high success rate across all simulation tasks. LeTO achieves a success rate comparable to diffusion policy and surpassing LSTM-GMM and IBC. Interestingly, when applying clipping to constrain the actions generated by the diffusion policy and LSTM-GMM, although this method sometimes increases the smoothness of the trajectory, it results in a reduction in their task success rates.
An intuitive explanation for this phenomenon is that both of these methods do not consider constraints during their training process; they merely fitted human demonstrations. When constraints are imposed on the actions generated by the trained models during deployment, it introduces additional factors that inevitably affect their performance. In contrast, our proposed LeTO, due to its ability to achieve end-to-end trajectory optimization while fitting demonstration data, balance the training objectives of satisfying constraints, smoothing the trajectories, and minimizing errors with demonstrations.
V Realworld Evaluation
For real robot experiments, we tackled constraints-critical tasks that demand smooth and safe trajectories, which are Move-the-stack (Fig. 5) and Arrange-chopsticks (Fig. 7). We use SpaceMouse teleoperating the end-effector of the robot manipulator to collect human demonstration data. More details of experiments and results can be seen in the attached video.
V-A Move-the-stack
[…]
Figure 5: Move-the-stack task. The robot grasps a set of stacked objects, smoothly transport them, and ultimately place them onto a black board, ensuring none of the stacked objects fall off. The black tape is for marking a consistent grasping position.
[…]
Figure 6: When chopsticks fall, which is a type of OOD data, both the diffusion policy and LeTO can still move the stacked cups and place them on the black board. In contrast, LSTM-GMM often exhibits erroneous behaviors, such as moving to incorrect locations or acting erratically.
TABLE VI: Metrics for evaluating the Move-the-stack task. “acc-peak” represents the highest acceleration achieved in any of the test trials while “avg-acc-mean/max/std” are the average of mean/max/std values across all test trials.
avg-acc-mean/max/std succ-rate acc-peak
LSTM-GMM 0.053/0.480/0.081 4/20 0.647
DiffusionPolicy 0.037/0.336/0.055 4/20 0.721
LeTO 0.037/0.219/0.042 13/20 0.250
[…]
Figure 7: Arrange-chopsticks task. The robot picks up randomly placed two chopsticks and arrange them neatly on a bowl.
TABLE VII: Metrics for evaluating the Arrange-chopsticks task. “contact-peak” represents the highest contact force achieved in any of the test trials while “avg-contact-max” is the average of max contact forces across all test trials.
avg-contact-max contact-peak succ-rate
LSTM-GMM 13.2 N 27.3 N 0/20
DiffusionPolicy 13.9 N 24.8 N 18/20
LeTO 11.7 N 16.1 N 17/20
The objective of the “Move-the-stack” task is that the robot grasps a set of stacked objects, smoothly transport them, and ultimately place them onto a black board, ensuring none of the stacked objects fall off. The initial position of the robot and the black board are randomized. The policy inputs end-effector positions and images from a third-person perspective camera, and controls the end-effector position. We introduce velocity constraint [-1,1] and acceleration constraint [-0.25,0.25] for LeTO. Notably, while we did not place chopsticks on top during human demonstration collection, a chopstick was added atop the stacked cups during policy rollout.
For imitation learning algorithms, the task is challenging because: 1) The policy needs to generate very smooth trajectories. Any roughness in the trajectory can lead to vibrations at the robot’s end-effector, causing the cups to shake. This vibration can accumulate and lead to the chopstick dropping. 2) Since a chopstick is introduced during policy rollout, this out-of-distribution (OOD) observation adds greater uncertainty to the network’s input. This uncertainty may make the network’s output jerkier, resulting in a poorer quality trajectory.
The results are shown in Table VI. From the results, LeTO excels in terms of trajectory smoothness and, correspondingly, also boasts the highest success rate. This demonstrates that due to the presence of end-to-end trajectory optimization, LeTO can robustly maintain its effectiveness even when facing OOD data. Another point worth noting is that when chopsticks fall (which is also a type of OOD data), both the diffusion policy and LeTO can still move the stacked cups and place them on the black board. In contrast, LSTM-GMM often exhibits erroneous behaviors, such as moving to incorrect locations or acting erratically (see Fig. 6 and the attached video). This showcases the robustness of LeTO and the diffusion policy.
V-B Arrange-chopsticks
The objective of the “Arrange-chopsticks” task is to pick up randomly placed two chopsticks and arrange them neatly on a bowl. The policy inputs end-effector poses and images from two cameras, a third-person perspective camera and a wrist-mounted camera, and controls the end-effector position, yaw angle, and grasping.
The challenges of this task include: 1. It is a relative long-horizon task, requiring the policy to have the capability to represent complex tasks. 2. Due to the slender nature of chopsticks, there is a high risk of the robot making forceful contact with the table when attempting to grasp them. For the sake of hardware safety, such collisions should be minimized as much as possible.
We introduce both position, velocity, and acceleration constraints for LeTO. During the demonstration, the lowest position of the robot’s end-effector on the z-axis was recorded at 0.0501 m. We have opted for 0.0475 m as the z-axis positional constraint within the LeTO framework. This choice is made with the task’s completion in mind, where the goal is not to prevent any contact between the robot’s end-effector and the table surface, but rather to allow for the slightest possible contact that still enables the picking up of slender chopsticks.
The results are shown in Table VII. Both LeTO and diffusion policy have high success rate, but the contact forces in LeTO tests are significantly smaller.
VI Discussion and future work
In this paper, we present LeTO, a framework for learning constrained visuomotor policy with differentiable trajectory optimization. By balancing the objective of minimizing errors with the need to satisfy trajectory constraints, LeTO allows for deployment in tasks where safety and reliability are paramount.
VI-A Hyperparameter, Training Time, and Training Stability
TABLE VIII: The training time for 20 epochs on the square ph task with a batch size of 64, conducted on a single GeForce RTX 3080. The hyperparameters for diffusion policy and LSTM-GMM are adopted from the default parameters reported in [12, 2], respectively. For the hyperparameters of LeTO, please refer to our open-source code.
Policy LSTM-GMM DiffusionPolicy LeTO
Training Time 2.90 hours 0.92 hours 7.25 hours
Regarding LeTO, its core neural network comprises CNN+LSTM, which is identical to the architecture of the basic imitation learning method BC-RNN. Therefore, we believe the discussion of hyperparameters of LeTO’s core network can refer to the discussion for BC-RNN in [12].
Concerning the hyperparameters of the DTO layer, since it comes from model-based formulation, it is possible to intuitively summarize how to select effective hyperparameters:
1. As shown in Remark 1, the optimization problem of the DTO layer is always feasible during training, given that the positions (and orientations) in the dataset meet the set position constraints. The consistent feasibility of the DTO layer ensures training stability, which is evident from the training logs of LeTO available in our open-source code.
2. In terms of constraints for acceleration and velocity, it is important to avoid overly small constraints. This could lead to a significant discrepancy between the generated trajectory and the dataset’s trajectory, adversely affecting model performance even with stable training. In our simulation experiments, LeTO trained at a maximum acceleration of 0.1 (normalized value) shows high performance, but setting it to 0.01 (normalized value) may result in worse outcomes. By analogy, the same consideration applies to the smoothing weight.
Table VIII compares the training times of LSTM-GMM, diffusion policy, and LeTO. We acknowledge that one limitation of LeTO is the computational overhead introduced by the differentiable optimization layer. Future work could look into improving the computational efficiency of the differentiable optimization process, potentially through the use of more efficient solvers or by learning approximations of the optimization layer.
VI-B LeTO in Low-data and High-data Regimes
We would like to highlight that networks with the training objective of minimizing the errors for fitting human demonstrated actions can readily incorporate our proposed DTO layer at the head. This suggests that for scalability, one could consider designing a LeTO variant with transformer, enhancing the model’s capability to scale to large datasets. However, the time and cost involved in training could still be a limiting factor. Therefore, future research focused on accelerating differentiable optimization is pivotal for scalability.
Furthermore, exploring policies that excel in low-data regimes based on differentiable optimization layers is an intriguing research direction. By constructing differentiable optimization layers using prior knowledge of the model, policy may reduce reliance on data.
VI-C LeTO in Reinforcement Learning
Due to the differentiability, the DTO layer can be designed to integrated into the robot learning pipelines, enabling end-to-end optimization and inference. This opens up the possibility of extending LeTO to reinforcement learning, offering a promising future research direction. By integrating differentiable trajectory optimization with reinforcement learning, it is possible to ensure that the generated actions comply with the designed constraints, thereby significantly enhancing the safety and interpretability of the policy.
VI-D LeTO in High-safety Scenarios
The safe and constraint-controlled output form of LeTO holds significant potential in robot learning scenarios where safety is of utmost importance, such as in human-robot interaction and surgical robotics. Investigating the application of LeTO, or more broadly, robot policies that integrate differential optimization in these settings, represents a highly interesting and promising research direction.